OpenStudy (anonymous):

Solve the equation for y and then find the value of y for the given x-value. 5x + y = 10 A. x = -2 B. x = 3

OpenStudy (austinl):

Ok, what do we do to cancel out an addition?

OpenStudy (anonymous):

subtract

OpenStudy (anonymous):

?

OpenStudy (austinl):

Correctomundo! So, what would we subtract to get y by iteslf?

OpenStudy (anonymous):

5?

OpenStudy (austinl):

That 5 is attached to a variable so you have to move that entire unit. What would it look like after you have done this?

OpenStudy (anonymous):

y=10

OpenStudy (austinl):

Nope, you have to move the 5 AND the variable it is attached to. \(\color{red}{5x} + \color{blue}{y} = \color{green}{10} \) So what do you think now?

OpenStudy (anonymous):

im not sureeee

OpenStudy (anonymous):

?

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

@austinL

OpenStudy (austinl):

Sorry, got distracted! You need to subtract 5x from both sides. \(\color{blue}{y} = \color{green}{10}-\color{red}{5x}\)

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

y = 5x?

OpenStudy (austinl):

No, What I put up above is what y is.

OpenStudy (anonymous):

oh so thats the answer

OpenStudy (austinl):

No, then according to the instructions you need to input those values of x and solve for y.

OpenStudy (anonymous):

then what?

OpenStudy (austinl):

That is all you need to do!

OpenStudy (anonymous):

input them wehre im so confused

OpenStudy (anonymous):

????????????????

OpenStudy (austinl):

Put them in where it says x.....

OpenStudy (anonymous):

but theres only one place to put one of the x's

OpenStudy (austinl):

Correct, it gives you two parts of the question. A. x = -2 B. x = 3 You then plug those into the equation that we made. y=10-5(x) Do you think you can solve it from there?

OpenStudy (anonymous):

can u show me cause im not getting what u saying

OpenStudy (austinl):

For part A, \(x=2\) We put that into the equation, \(y=10-5x\Rightarrow10-5(2)\) \(y=0\) Make sense? Now try part B for yourself! :D

OpenStudy (anonymous):

it was -2

OpenStudy (austinl):

Ok, then \(y=20\)